OpenStudy (anonymous):

(3x -14)/2x = 3x - 2 I have absolutely no idea how to do this. Can you please help me understand the steps I need to take to solve this?

OpenStudy (kropot72):

First multiply both sides by 2x. This will get rid of the fraction. Can you do this?

OpenStudy (anonymous):

Yes, now I have 3x - 14 = 6x^2 - 4x

OpenStudy (kropot72):

Next take all the terms on the left hand side to the right hand side (remembering to change signs) to get 6x^2...................................... = 0

OpenStudy (anonymous):

So I add 3x, then subtract 14 which leaves me with 0 = 6x^2 - 7x - 14

OpenStudy (kropot72):

The rearrangement of terms might be clearer shown as follows: 3x - 14 -3x + 14 = 6x^2 - 4x -3x + 14 0 = 6x^2 - 7x + 14

OpenStudy (anonymous):

\[\frac{ 3x-14 }{ 2x } = 3x - 2\]

OpenStudy (anonymous):

Cross Multiply :\[3x-14 = 2x ( 3x-2)\]

OpenStudy (anonymous):

\[3x - 14 = 6x^2 - 4x\]

OpenStudy (anonymous):

Collect Like terms : \[6x^2 - 4x-3x + 14 =0\] \[6x^2 - 7x + 14 = 0\]

OpenStudy (anonymous):

Nw it is a Quadratic Eq use Quad Formula: \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Hope this Helps:

OpenStudy (anonymous):

@JamesR4494 Any Doubts: ?

OpenStudy (kropot72):

The next step is to solve the quadratic 6x^2 - 7x + 14 = 0 using the formula: \[x=\frac{-b \pm \sqrt{(b ^{2}}-4ac)}{2a}\]

OpenStudy (kropot72):

a = 6 b = -7 c = 14

OpenStudy (anonymous):

I understand everything you guys previously wrote but I don't know how to solve this problem using the quadratic formula, I don't quite understand how to apply it to this problem.

OpenStudy (anonymous):

If u Dont knw this Just FActorize the Eq

OpenStudy (anonymous):

6x^2 - 7x + 14 Product = 84 Sum = -7 Find Two such Numbers: thhis method..Sucks...)

OpenStudy (anonymous):

x+ y = -7 xy =84

OpenStudy (anonymous):

Oh my... I made this equation much harder than it actually is... I apologize. I wrote it incorrectly. The real equation is (3x - 14)/-2 = 3x - 2 This explains alot...

OpenStudy (anonymous):

I truly apologize for wasting both of your guy's time. I have solved the equation. Thanks for the help even if it wasn't the right equation it's nice to know that if I'm ever stumped on a question there are helpful members willing to guide me to the solution.